# In Search of a Hidden Curve

### v1

@@ 2.2. Simple Hurwitz numbers [51]
Recently there have been totally unexpected spectacular developments in the topology of moduli spaces ${\mathcal{M}}_{g},{\mathcal{M}}_{g,n}$, and ¯${\overline{\mathcal{M}}}_{g,n}$ (see for example, [16]). From the point of view of In Search of a Hidden Curve, and from the perspective of Norbury discovering the Norbury classes in $H^{*}({\overline{\mathcal{M}}}_{g,n},{\mathbb{Q}})$ using topological recursion [82], which were later identified as the Euler classes associated with a supersymmetric generalization of the work of Mirzakhani by Stanford-Witten [90], we wonder if there could be a spectral curve behind the recent developments.
[…]
@@
is a topological covering of degree $d$. Each $z_{k}\in B$ is a branched point of $h$, and a critical point (i.e., a zero of $dh$) on $C$ is called a ramification point of $h$. Since $dh$ has only simple zeros with distinct critical values of $h$, the number of ramification points of $h$, except for the poles, is equal to the number of branched points, which we denote by $r$. Therefore, $h^{-1}(z_{k})$ consists of $d-1$ points. Then by comparing the Euler characteristic of the covering space and its base space, we obtain
@@
  • $\mu(\hat{i})$ is the vector of $\ell-1$ entries obtained by deleting the $i$-th entry $\mu_{i}$.
  • $(\mu(\hat{i}),\alpha,\beta)$ is the vector of $\ell+1$ entries obtained by appending two new entries $\alpha$ and $\beta$ to $\mu(\hat{i})$.
  • $\mu(\hat{i},\hat{j})$ is the vector of $\ell-2$ entries obtained by deleting the $i$-th and the $j$-th entries $\mu_{i}$ and $\mu_{j}$.
  • $(\mu(\hat{i},\hat{j}),\mu_{i}+\mu_{j})$ is the vector of $\ell-1$ entries obtained by appending a new entry $\mu_{i}+\mu_{j}$ to $\mu(\hat{i},\hat{j})$.
The final sum is over all partitions of $g$ into non-negative integers $g_{1}$ and $g_{2}$, and a disjoint union decomposition (or a set partition) of entries of $\mu(\hat{i})$ as a set, allowing the empty set.
@@
where $\widetilde{X}$ is the universal covering space of ${\mathbb{P}}^{1}\setminus B$, and $[d]=\{1,2,\dots,d\}$ is the index set on which $\pi_{1}({\mathbb{P}}^{1}\setminus B)$ acts via the representation $\rho$.
To make $X_{\rho}$ a Hurwitz covering, we need to specify the monodromy of the representation at each branch point of $B$. Let $\{\gamma_{1},\dots,\gamma_{r},\gamma_{\infty}\}$ denote the collection of non-intersecting loops on ${\mathbb{P}}^{1}$, starting from $0\in{\mathbb{P}}^{1}$, rotating around $z_{k}$ counter-clockwise, and coming back to 00, for each $k=1,\dots,r$. The last loop $\gamma_{\infty}$ does the same for $\infty\in{\mathbb{P}}^{1}$. Since ${\mathbb{P}}^{1}$ is simply connected, we have
@@
2.3. Fugue in two themes: The Laplace transform of the simple Hurwitz numbers
Let us now compute the Laplace transform of the simple Hurwitz number $H_{g}(\mu)$, considered as a function in $\mu\in{\mathbb{Z}}_{+}^{\ell}$. According to Principle 2.3, the result should give us the mirror dual of simple Hurwitz numbers. We explain where the Lambert curve $x=ye^{1-y}$ comes from, and its essential role in computing the Laplace transform. The most surprising feature is that the result of the Laplace transform of $H_{g}(\mu)$ is a polynomial if $02g-2+\ell>$ 0. This polynomiality produces powerful consequences, which is the main subject of the following lectures.
The most important reason why we are interested in Hurwitz numbers in this summer school on complex Lagrangians is because of the theorem due to Ekedahl, Lando, Shapiro and Vainshtein [35] that relates the simple Hurwitz numbers with the intersection numbers of tautological classes on the moduli spaces of curves. From analyzing their formula, we see the emergence of the spectral curve, which we consider as a complex Lagrangian in the holomorphic symplectic surface.
Let us recall the necessary notations here. Our main object is the moduli stack ¯${\overline{\mathcal{M}}}_{g,\ell}$ consisting of stable algebraic curves of genus $g\geq 0$ with $\ell\geq 1$ distinct smooth labeled points. Forgetting the last labeled point on a curve gives a canonical projection
@@ 2.3. Fugue in two themes: The Laplace transform of the simple Hurwitz numbers
Since the last labeled point moves on the curve, the projection $\pi$ can be considered as a universal family of $\ell$-pointed curves. This is because for each point $[C,(x_{1},\dots,x_{\ell})]\in{\overline{\mathcal{M}}}_{g,\ell}$, the fiber of $\pi$ is indeed $C$ itself.
If $x_{\ell+1}\in C$ is a smooth point of $C$ other than $\{x_{1},\dots,x_{\ell}\}$, then this point represents an element $[C,(x_{1},\dots,x_{\ell+1})]\in{\overline{\mathcal{M}}}_{g,\ell+1}$. If $x_{\ell+1}=x_{i}$ for some $i=1,\dots,\ell$, then this point represents a stable curve obtained by attaching a rational curve ${\mathbb{P}}^{1}$ to $C$ at the original location of $x_{i}$, while carrying three special points on it. One is the singular point at which $C$ and ${\mathbb{P}}^{1}$ intersects. The other two points are labeled as $x_{i}$ and $x_{\ell+1}$. And if $x_{\ell+1}\in C$ coincides with one of the nodal points of $C$, say $x\in C$, then this point represents another stable curve. This time, consider the local normalization $\widetilde{C}\rightarrow C$ about the singular point $x\in C$, and let $x_{+}$ and $x_{-}$ be the two points in the fiber. The stable curve we have is the curve $\widetilde{C}\cup{\mathbb{P}}^{1}$, where the two curves intersect at $x_{+}$ and $x_{-}$. The labeled point $x_{\ell+1}$ is placed on the attached ${\mathbb{P}}^{1}$ different from these two singular points. Because ${\mathbb{P}}^{1}$ with three distinct labeled points form a moduli space consisting of a single point, the processes above (called stabilization) leave no ambiguity.
@@
A universal family produces what we call the tautological bundles on the moduli space. The cotangent sheaf $T^{*}C=\omega_{C}$ of each fiber of $\pi$ is glued together to form a relative dualizing sheaf $\omega$ on ¯${\overline{\mathcal{M}}}_{g,\ell+1}$. The push-forward
@@
is such a tautological vector bundle on ¯${\overline{\mathcal{M}}}_{g,\ell}$ of fiber dimension
@@
and is called the Hodge bundle on ¯${\overline{\mathcal{M}}}_{g,\ell}$. By assigning $x_{\ell+1}=x_{i}$ to each $[C,(x_{1},\dots,x_{\ell})]\in{\overline{\mathcal{M}}}_{g,\ell}$, we construct a section
@@
on ¯${\overline{\mathcal{M}}}_{g,\ell}$. The fiber of ${\mathbb{L}}_{i}$ at $[C,(x_{1},\dots,x_{\ell})]$ is identified with the cotangent line $T_{x_{i}}^{*}C$.
The tautological classes of ¯${\overline{\mathcal{M}}}_{g,\ell}$ are rational cohomology classes including
$\psi_{i}=c_{1}({\mathbb{L}}_{i})\in H^{2}({\overline{\mathcal{M}}}_{g,\ell},{\mathbb{Q}})\quad\text{and}\quad\lambda_{j}=c_{j}({\mathbb{E}})\in H^{2j}({\overline{\mathcal{M}}}_{g,\ell},{\mathbb{Q}}).$
@@
The simple Hurwitz numbers are expressible as the intersection numbers of tautological classes on the moduli space ¯${\overline{\mathcal{M}}}_{g,\ell}$ as follows. Let $\mu\in{\mathbb{Z}}_{+}^{\ell}$ be a positive integer vector. Then we have
@@
It is not our purpose to give a proof of the ELSV formula in these lectures. There are excellent articles by now about this remarkable formula. We refer to [60, 85].
@@
(2.10) $\frac{k^{k+n}}{k!}e^{-k}\sim\frac{1}{\sqrt{2\pi}}\;k^{n-{\frac{1}{2}}},\qquad k\gg 0$
@@
[…]
@@
Note that for every $n>0$, the summation in the definition of $\xi_{n}(w)$ in (2.11) can be taken from $k=0$ to $\infty$. For $n=0$, the $k=0$ term contributes $10^{0}=$ 1 in the summation. So let us define
@@
(2.19) $-\frac{d}{dw}=x\frac{d}{dx}=\frac{y}{1-y}\frac{d}{dy}.$
@@
(2.21) $-\frac{d}{dw}=x\frac{d}{dx}=\frac{y}{1-y}\frac{d}{dy}=t^{2}(t-1)\frac{d}{dt}.$
@@
For every $n\geq 0$, ${\hat{\xi}}_{n}(t)$ is a polynomial in $t$ of degree $12n+$ 1. For $n>0$ it has an expansion
@@
We note that there is no a priori reason for the Laplace transform of $H_{g}(\mu)$ to be a polynomial. Because it is a polynomial, we obtain a polynomial generating function of linear Hodge integrals ${\langle}\tau_{n_{1}}\cdots\tau_{n_{\ell}}\lambda_{j}{\rangle}_{g,\ell}$. We utilize this polynomiality in the concluding Fuge below.
@@
The motivation of the author to work on Catalan numbers [33, 31] was to find an analytically simpler example of enumeration problem for which a similar polynomiality holds. The coordinate change for the Catalan case was obtained from looking for an analogy of (2.20).
@@
2.4. Fugue in three themes: Two quantum curves for the Lambert curve
@@
The difference is only in the constant multiplication $x\longmapsto ex$. For the calculation of quantum curves, we use the classical one. This works better for quantum curves, because the differential operator we need is $D:=x\frac{d}{dx}$, which is invariant under the ${\mathbb{C}}^{*}$-action.
@@
$y=\frac{t-1}{t}\quad\text{and}\quad x=ye^{-y}.$
@@
    It restricts to the top degree terms, and recovers the Dijkgraaf-Verlinde-Verlinde formula, or the Virasoro constraint condition, for the $\psi$-class intersection numbers on ¯${\overline{\mathcal{M}}}_{g,\ell}$ [95].
@@
where $D_{i}=t_{i}^{2}(t_{i}-1)\frac{\partial}{\partial t_{i}}$. As before, []=$[\ell]=\{1,\dots,\ell\}$ is the index set, and $[\ell;\hat{i}]$ is the index set obtained by deleting $i$ from $[\ell]$. The last summation is taken over all partitions $g=g_{1}+g_{2}$ of the genus $g$ and disjoint union decompositions $J\sqcup K=[\ell;\hat{i}]$ satisfying the stability conditions $02g_{1}-1+|J|>$ 0 and $02g_{2}-1+|K|>$ 0. For a subset $I\subset[\ell]$ we write $t_{I}=(t_{i})_{i\in I}$.
@@
(2.33) $\sum_{n_{[\ell]}}{\langle}\tau_{n_{[\ell]}}\Lambda_{g}^{\vee}(1){\rangle}_{g,\ell}\left((2g-2+\ell){\hat{\xi}}_{n_{[\ell]}}(t_{[\ell]})+\sum_{i=1}^{\ell}\frac{1}{t_{i}}\;{\hat{\xi}}_{n_{i}+1}(t_{i}){\hat{\xi}}_{[\ell;\hat{i}]}(t_{[\ell;\hat{i}]})\right)\\ =\sum_{i<j}\sum_{m,n_{[\ell;\hat{i}\hat{j}]}}{\langle}\tau_{m}\tau_{n_{[\ell;\hat{i}\hat{j}]}}\Lambda_{g}^{\vee}(1){\rangle}_{g,\ell-1}{\hat{\xi}}_{n_{[\ell;\hat{i}\hat{j}]}}(t_{[\ell;\hat{i}\hat{j}]})\frac{{\hat{\xi}}_{m+1}(t_{i}){\hat{\xi}}_{0}(t_{j})t_{i}^{2}-{\hat{\xi}}_{m+1}(t_{j}){\hat{\xi}}_{0}(t_{i})t_{j}^{2}}{t_{i}-t_{j}}\\ +{\frac{1}{2}}\sum_{i=1}^{\ell}\sum_{n_{[\ell;\hat{i}]}}\sum_{a,b}\Bigg{(}{\langle}\tau_{a}\tau_{b}\tau_{n_{[\ell;\hat{i}]}}\Lambda_{g-1}^{\vee}(1){\rangle}_{g-1,\ell+1}\\ +\sum_{\begin{subarray}{c}g_{1}+g_{2}=g\\ I\coprod J=[\ell;\hat{i}]\end{subarray}}^{\rm{stable}}{\langle}\tau_{a}\tau_{n_{I}}\Lambda_{g_{1}}^{\vee}(1){\rangle}_{g_{1},|I|+1}{\langle}\tau_{b}\tau_{n_{J}}\Lambda_{g_{2}}^{\vee}(1){\rangle}_{g_{2},|J|+1}\Bigg{)}{\hat{\xi}}_{a+1}(t_{i}){\hat{\xi}}_{b+1}(t_{i}){\hat{\xi}}_{n_{[\ell;\hat{i}]}}(t_{[\ell;\hat{i}]}),$
where []=$[\ell]=\{1,2\dots,\ell\}$ is the index set, and for a subset $I\subset[\ell]$, we denote
$t_{I}=(t_{i})_{i\in I},\quad n_{I}=\{\,n_{i}\,|\,i\in I\,\},\quad\tau_{n_{I}}=\prod_{i\in I}\tau_{n_{i}},\quad{\hat{\xi}}_{n_{I}}(t_{I})=\prod_{i\in I}{\hat{\xi}}_{n_{i}}(t_{i}).$
@@
The Deligne-Mumford stack ¯$\overline{\mathcal{M}}_{g,\ell}$ is defined as the moduli space of stable curves satisfying the stability condition $02-2g-\ell<$ 0. However, Hurwitz numbers are well defined for unstable geometries $(g,\ell)=(0,1)$ and $(0,2)$. It is an elementary exercise (of tree counting, see [62]) to show that
@@
(2.35) ¯$\int_{\overline{{\mathcal{M}}}_{0,1}}\frac{1}{1-k\psi}=\frac{1}{k^{2}},$
(2.36) ¯$\int_{\overline{{\mathcal{M}}}_{0,2}}\frac{1}{(1-\mu_{1}\psi_{1})(1-\mu_{2}\psi_{2})}=\frac{1}{\mu_{1}+\mu_{2}}.$
@@
2.5. Fugue in four themes: New proofs of Witten-Kontsevich and the $\lambda_{g}$ formulas
Now let us move to proving the Witten conjecture and the $\lambda_{g}$-formula using our recursion (2.33). Although these important formulas have been proved a long time ago, we present a new proofs here which are much simpler, just to illustrate the power of the topological recursion type formula.

### v2

@@ 2.2. Simple Hurwitz numbers
Recently there have been totally unexpected spectacular developments in the topology of moduli spaces ${\mathcal{M}}_{g},{\mathcal{M}}_{g,n}$, and ${\overline{\mathcal{M}}}_{g,n}$ (see for example, [17]). From the point of view of In Search of a Hidden Curve, and from the perspective of Norbury discovering the Norbury classes in $H^{*}({\overline{\mathcal{M}}}_{g,n},{\mathbb{Q}})$ using topological recursion [85], which were later identified as the Euler classes associated with a supersymmetric generalization of the work of Mirzakhani by Stanford-Witten [93], we wonder if there could be a spectral curve behind the recent developments.
A simple Hurwitz number represents the number of a particular type of meromorphic functions defined on an algebraic curve $C$ of genus $g$. Let $\mu=(\mu_{1},\dots,\mu_{\ell})\in{\mathbb{Z}}_{+}^{\ell}$ be a partition of a positive integer $d$ of length $\ell$. This means that $|\mu|\overset{\text{def}}{=}\mu_{1}+\cdots+\mu_{\ell}=d$. Instead of ordering parts of $\mu$ in the decreasing order, we consider them as a vector consisting of $\ell$ positive integers. By a Hurwitz covering of type $(g,\mu)$ we mean a meromorphic function $h:C\rightarrow{\mathbb{C}}$ that has $\ell$ labelled poles $\{x_{i},\dots,x_{\ell}\}$, such that the pole order of $h$ at $x_{i}$ is $\mu_{i}$ for every $i=1,\dots,\ell$, and that except for these poles, the holomorphic $1$-form dhdh has simple zeros on $C\setminus\{x_{i},\dots,x_{\ell}\}$ with distinct critical values of $h$. A meromorphic function of $C$ is a holomorphic map of $C$ onto ${\mathbb{P}}^{1}$. In algebraic geometry, the situation described above is summarized as follows: $h:C\rightarrow{\mathbb{P}}^{1}$ is a ramified covering of ${\mathbb{P}}^{1}$, simply ramified except for $\infty\in{\mathbb{P}}^{1}$. We identify two Hurwitz coverings $h_{1}:C_{1}\rightarrow{\mathbb{P}}^{1}$ with poles at $\{x_{1},\dots,x_{\ell}\}$ and $h_{2}:C_{2}\rightarrow{\mathbb{P}}^{1}$ with poles at $\{y_{1},\dots,y_{\ell}\}$ if there is a biholomorphic map $\phi:C_{1}\overset{\sim}{\rightarrow}C_{2}$ such that $\phi(x_{i})=y_{i}$, $i=1,\dots,\ell$, and
${C_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h_{1}}$$\scriptstyle{\phi}\scriptstyle{\sim}$${C_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h_{2}}$${{\mathbb{P}}^{1}}$.
@@
is a topological covering of degree $d$. Each $z_{k}\in B$ is a branched point of $h$, and a critical point (i.e., a zero of $dh$) on $C$ is called a ramification point of $h$. Since $dh$ has only simple zeros with distinct critical values of $h$, the number of ramification points of $h$, except for the poles, is equal to the number of branched points, which we denote by $r$. Therefore, $h^{-1}(z_{k})$ consists of d1d-1 points. Then by comparing the Euler characteristic of the covering space and its base space, we obtain
@@
[…]
@@
where $\widetilde{X}$ is the universal covering space of ${\mathbb{P}}^{1}\setminus B$, and $d]=\{1,2,\dots,d\}$ is the index set on which $\pi_{1}({\mathbb{P}}^{1}\setminus B)$ acts via the representation $\rho$.
To make $X_{\rho}$ a Hurwitz covering, we need to specify the monodromy of the representation at each branch point of $B$. Let $\{\gamma_{1},\dots,\gamma_{r},\gamma_{\infty}\}$ denote the collection of non-intersecting loops on ${\mathbb{P}}^{1}$, starting from $0\in{\mathbb{P}}^{1}$, rotating around $z_{k}$ counter-clockwise, and coming back to 0, for each $k=1,\dots,r$. The last loop $\gamma_{\infty}$ does the same for $\infty\in{\mathbb{P}}^{1}$. Since ${\mathbb{P}}^{1}$ is simply connected, we have
@@
2.3. The Laplace transform of the simple Hurwitz numbers
Let us now compute the Laplace transform of the simple Hurwitz number $H_{g}(\mu)$, considered as a function in $\mu\in{\mathbb{Z}}_{+}^{\ell}$. According to Principle 2.3, the result should give us the mirror dual of simple Hurwitz numbers. We explain where the Lambert curve $x=ye^{1-y}$ comes from, and its essential role in computing the Laplace transform. The most surprising feature is that the result of the Laplace transform of $H_{g}(\mu)$ is a polynomial if $2g-2+\ell>0$. This polynomiality produces powerful consequences, which is the main subject of the following subsections.
The most important reason for our interest in Hurwitz numbers in this summer school on complex Lagrangians lies in the theorem due to Ekedahl, Lando, Shapiro and Vainshtein [36] that relates the simple Hurwitz numbers with the intersection numbers of tautological classes on the moduli spaces of curves. From analyzing their formula, we see the emergence of the spectral curve, which we consider as a complex Lagrangian in the holomorphic symplectic surface.
Let us recall the necessary notations here. Our main object is the moduli stack ${\overline{\mathcal{M}}}_{g,\ell}$ consisting of stable algebraic curves of genus $g\geq 0$ with $\ell\geq 1$ distinct smooth labeled points. Forgetting the last labeled point on a curve gives a canonical projection
@@ 2.3. The Laplace transform of the simple Hurwitz numbers
Since the last labeled point moves on the curve, the projection $\pi$ can be considered as a universal family of $\ell$-pointed curves. This is because for each point $C,(x_{1},\dots,x_{\ell})]\in{\overline{\mathcal{M}}}_{g,\ell}$, the fiber of $\pi$ is indeed $C$ itself.
If $x_{\ell+1}\in C$ is a smooth point of $C$ other than $\{x_{1},\dots,x_{\ell}\}$, then this point represents an element $C,(x_{1},\dots,x_{\ell+1})]\in{\overline{\mathcal{M}}}_{g,\ell+1}$. If $x_{\ell+1}=x_{i}$ for some $i=1,\dots,\ell$, then this point represents a stable curve obtained by attaching a rational curve ${\mathbb{P}}^{1}$ to $C$ at the original location of $x_{i}$, while carrying three special points on it. One is the singular point at which $C$ and ${\mathbb{P}}^{1}$ intersects. The other two points are labeled as $x_{i}$ and $x_{\ell+1}$. And if $x_{\ell+1}\in C$ coincides with one of the nodal points of $C$, say $x\in C$, then this point represents another stable curve. This time, consider the local normalization $\widetilde{C}\rightarrow C$ about the singular point $x\in C$, and let $x_{+}$ and $x_{-}$ be the two points in the fiber. The stable curve we have is the curve $\widetilde{C}\cup{\mathbb{P}}^{1}$, where the two curves intersect at $x_{+}$ and $x_{-}$. The labeled point $x_{\ell+1}$ is placed on the attached ${\mathbb{P}}^{1}$ different from these two singular points. Because ${\mathbb{P}}^{1}$ with three distinct labeled points form a moduli space consisting of a single point, the processes above (called stabilization) leave no ambiguity.
@@
A universal family produces what we call the tautological bundles on the moduli space. The cotangent sheaf $T^{*}C=\omega_{C}$ of each fiber of $\pi$ is glued together to form a relative dualizing sheaf $\omega$ on ${\overline{\mathcal{M}}}_{g,\ell+1}$. The push-forward
@@
is such a tautological vector bundle on ${\overline{\mathcal{M}}}_{g,\ell}$ of fiber dimension
@@
and is called the Hodge bundle on ${\overline{\mathcal{M}}}_{g,\ell}$. By assigning $x_{\ell+1}=x_{i}$ to each $C,(x_{1},\dots,x_{\ell})]\in{\overline{\mathcal{M}}}_{g,\ell}$, we construct a section
@@
on ${\overline{\mathcal{M}}}_{g,\ell}$. The fiber of ${\mathbb{L}}_{i}$ at $C,(x_{1},\dots,x_{\ell})]$ is identified with the cotangent line $T_{x_{i}}^{*}C$.
The tautological classes of ${\overline{\mathcal{M}}}_{g,\ell}$ are rational cohomology classes including
$\psi_{i}=c_{1}({\mathbb{L}}_{i})\in H^{2}({\overline{\mathcal{M}}}_{g,\ell},{\mathbb{Q}})\qquad\text{and}\qquad\lambda_{j}=c_{j}({\mathbb{E}})\in H^{2j}({\overline{\mathcal{M}}}_{g,\ell},{\mathbb{Q}}).$
@@
The simple Hurwitz numbers are expressible as the intersection numbers of tautological classes on the moduli space ${\overline{\mathcal{M}}}_{g,\ell}$ as follows. Let $\mu\in{\mathbb{Z}}_{+}^{\ell}$ be a positive integer vector. Then we have
@@
It is not our purpose to give a proof of the ELSV formula in these lectures. There are excellent articles about this remarkable formula. We refer to [64, 88].
@@
(2.10) $\frac{k^{k+n}}{k!}e^{-k}\sim\frac{1}{\sqrt{2\pi}}\;k^{n-{\frac{1}{2}}},\hskip 23.49976ptk\gg 0$
@@
$\xi_{n}(w)\sim\int_{0}^{\infty}\frac{1}{\sqrt{2\pi}}x^{n-{\frac{1}{2}}}e^{-xw}dx=\frac{\Gamma(n+{\frac{1}{2}})}{\sqrt{2\pi}\;w^{n+{\frac{1}{2}}}}.$
@@
Note that for every $n>0$, the summation in the definition of $\xi_{n}(w)$ in (2.11) can be taken from $k=0$ to $\infty$. For $n=0$, the $k=0$ term contributes $0^{0}=1$ in the summation. So let us define
@@
(2.19) $\frac{d}{dw}=x\frac{d}{dx}=\frac{y}{1-y}\frac{d}{dy}.$
@@
(2.21) $\frac{d}{dw}=x\frac{d}{dx}=\frac{y}{1-y}\frac{d}{dy}=t^{2}(t-1)\frac{d}{dt}.$
@@
For every $n\geq 0$, ${\hat{\xi}}_{n}(t)$ is a polynomial in $t$ of degree $2n+1$. For $n>0$ it has an expansion
@@
We note that there is no a priori reason for the Laplace transform of $H_{g}(\mu)$ to be a polynomial. Because it is a polynomial, we obtain a polynomial generating function of linear Hodge integrals ${\langle}\tau_{n_{1}}\cdots\tau_{n_{\ell}}\lambda_{j}{\rangle}_{g,\ell}$. We utilize this polynomiality in the concluding subsection below.
@@
The motivation for the author to work on Catalan numbers [34, 32] was to find an analytically simpler example of enumeration problem for which a similar polynomiality holds. The coordinate change for the Catalan case was obtained from looking for an analogy of (2.20).
@@
2.4. Two quantum curves for the Lambert curve
@@
[…]
@@
$y=\frac{t-1}{t}\qquad\text{and}\qquad x=ye^{-y}.$
@@
    It restricts to the top degree terms, and recovers the Dijkgraaf-Verlinde-Verlinde formula, or the Virasoro constraint condition, for the $\psi$-class intersection numbers on ${\overline{\mathcal{M}}}_{g,\ell}$ [98].
@@
where $D_{i}=t_{i}^{2}(t_{i}-1)\frac{\partial}{\partial t_{i}}$. As before, $\ell]=\{1,\dots,\ell\}$ is the index set, and $\ell;\hat{i}$] is the index set obtained by deleting $i$ from $\ell$]. The last summation is taken over all partitions $g=g_{1}+g_{2}$ of the genus $g$ and disjoint union decompositions $J\sqcup K=[\ell;\hat{i}$] satisfying the stability conditions $2g_{1}-1+|J|>0$ and $2g_{2}-1+|K|>0$. For a subset $I\subset[\ell]$ we write $t_{I}=(t_{i})_{i\in I}$.
@@
[…]
where $\ell]=\{1,2\dots,\ell\}$ is the index set, and for a subset $I\subset[\ell]$, we denote
$t_{I}=(t_{i})_{i\in I},\qquad n_{I}=\{\,n_{i}\,|\,i\in I\,\},\qquad\tau_{n_{I}}=\prod_{i\in I}\tau_{n_{i}},\qquad{\hat{\xi}}_{n_{I}}(t_{I})=\prod_{i\in I}{\hat{\xi}}_{n_{i}}(t_{i}).$
@@
The Deligne-Mumford stack $\overline{\mathcal{M}}_{g,\ell}$ is defined as the moduli space of stable curves satisfying the stability condition $2-2g-\ell<0$. However, Hurwitz numbers are well defined for unstable geometries $(g,\ell)=(0,1)$ and $(0,2)$. It is an elementary exercise (of tree counting, see [66]) to show that
@@
(2.35) $\int_{\overline{{\mathcal{M}}}_{0,1}}\frac{1}{1-k\psi}=\frac{1}{k^{2}},$
(2.36) $\int_{\overline{{\mathcal{M}}}_{0,2}}\frac{1}{(1-\mu_{1}\psi_{1})(1-\mu_{2}\psi_{2})}=\frac{1}{\mu_{1}+\mu_{2}}.$
@@
2.5. New proofs of Witten-Kontsevich and the $\lambda_{g}$ formulas
Now let us move to proving the Witten conjecture and the $\lambda_{g}$-formula using our recursion (2.33). Although these important formulas have been proved a long time ago, we present simpler proofs here just to illustrate the power of the topological recursion type formula.